Could someone please help me with this problem?
3x - 2(x-4)= 5x + 8

Found 2 solutions by ankor@dixie-net.com, Alan3354:
Answer by ankor@dixie-net.com(22740)

3x - 2(x-4)= 5x + 8
This is not going be solved, multiply what's inside the brackets (distribute)
3x - 2x + 8 = 5x + 8
subtract 8 from both sides
x = 5x + 8 - 8
x = 5x, obviously makes no sense,

Answer by Alan3354(69443)

Could someone please help me with this problem?
3x - 2(x-4)= 5x + 8
-----------
3x - 2x + 8 = 5x + 8
x = 5x
0 = 4x
x = 0